OpenStudy (kj4uts):

What is the average atomic weight (amu) of an element if the two known isotopes of the element have the abundances as shown below? Isotope 1: Mass = 31.163 amu; abundance = 35.16% Isotope 2: Mass = 34.296 amu; abundance = 64.84% A. 30.197 B. 35.221 C. 34.016 D. 33.194 Please explain. Thank you!

OpenStudy (kj4uts):

I know this is more of a chemistry problem but does anyone know how to solve this?

OpenStudy (kj4uts):

Do I add the numbers and divide by two?

OpenStudy (mathmate):

For more chemistry problems, please post in the chemistry section. If there is no response after a while, post the \(link\) here for further help. This way you get your answer faster. What you need to do is called the weighted average, namely the sum of the atomic weight multiplied by the abundance. The sum of abundance must equal 1 (or 100%). Here what you do is therefore: Isotope 1: 31.163 amu \(\times\) 35.16% =10.957 Isotope 2: 34.296 amu; \(\times\) 64.84%=22.238 Sum: = 10.957+22.238=33.194 rounded to 33.19 Note: In chemistry and physics, remember to round the answer to the proper number of significant figures. For multiplications, the number is the least significant of the multiplicants, namely four significant figures.

OpenStudy (kj4uts):

Thank you very much for your time and help! I appreciate that you explained the steps so that I can apply that to my other problems and next time I will post in the chemistry section :)

OpenStudy (mathmate):

You're welcome! :)